# Non-Intrusive Graph-Based Bot Detection for E-Commerce Using Inductive Graph Neural Networks

### v3

@@ Abstract
Malicious bots increasingly abuse e-commerce services (scraping, inventory hoarding, fraud) while evading conventional defenses. IP/rule blocking is brittle under proxies and botnets, and CAPTCHAs add friction yet are increasingly bypassed. We propose a non-intrusive, deployment-friendly framework that models session–URL interactions as a bipartite graph and applies an inductive GNN (GraphSAGE) to classify session nodes. By combining topology with lightweight behavioral and URL semantics, the model detects “feature-normal” automation. On real-world traffic with high-confidence bot labels, GraphSAGE outperforms a strong session-feature MLP baseline in AUC and F1. We further test robustness under adversarial edge perturbations and inductive generalization in a cold-start simulation; performance degrades modestly under mild perturbation and remains strong on unseen nodes, supporting real-time deployment without client-side instrumentation.
@@ I Introduction
E-commerce platforms face persistent automated abuse (scraping, inventory hoarding, fraud) from bots that increasingly mimic human browsing. Common defenses are either brittle (IP/user-agent blocking under proxy rotation) or intrusive (CAPTCHAs), motivating passive, high-precision detection from backend telemetry.
Graphs compactly capture relationships across sessions and content, enabling detection beyond per-session aggregates and improving robustness to evolving behavior (e.g., BotChase [1]). We construct a bipartite session–URL interaction graph from standard logs and apply an inductive GNN, GraphSAGE [2], to classify session nodes. Inductive learning supports cold-start scoring for new sessions and new pages; we ask whether this non-intrusive graph signal improves accuracy while remaining robust to graph perturbations and temporal shift.
@@
  • Non-intrusive graph formulation: We cast bot detection as learning on a session–URL graph built from standard server logs, avoiding CAPTCHAs and client-side instrumentation.
  • Inductive GraphSAGE with lightweight features: We use GraphSAGE to aggregate session and URL attributes, enabling immediate scoring of unseen session/URL nodes via learned neighborhood aggregation.
  • Robustness + deployability evaluation: We report gains over a session-feature MLP baseline and study adversarial edge perturbations and cold-start generalization, showing modest degradation under moderate perturbation and strong performance on unseen nodes.
@@ VI Conclusion
We presented a non-intrusive, graph-based bot detection framework evaluated on real-world traffic with high-confidence bot labels that models session–URL interactions and applies inductive GraphSAGE for session classification. By combining topology with semantic features, the refined-graph model improves over a strong session-only MLP while avoiding CAPTCHAs and client-side instrumentation. Robustness experiments show modest degradation under edge perturbations, and cold-start evaluation confirms strong inductive generalization to new sessions/URLs, supporting deployment as a bounded-latency backend scoring module.

### v4

@@ Abstract
Malicious bots abuse e-commerce services while evading conventional defenses. IP/rule blocking is brittle under proxy rotation, and CAPTCHAs add friction yet are often bypassed. We propose a non-intrusive framework that models session–URL interactions as a bipartite graph and uses an inductive GNN (GraphSAGE) to classify session nodes. Combining topology with lightweight behavioral and URL semantics enables detection of “feature-normal” automation. On real-world traffic with high-confidence bot labels, GraphSAGE outperforms a session-feature MLP baseline in AUC and F1, and remains robust under mild adversarial edge perturbations and in cold-start inductive evaluation—supporting real-time deployment without client-side instrumentation.
@@ I Introduction
E-commerce platforms face persistent automated abuse from bots that mimic human browsing. Common defenses are brittle under proxy rotation or intrusive (CAPTCHAs), motivating passive detection from backend telemetry.
Graphs compactly capture relationships across sessions and content, enabling detection beyond per-session aggregates (e.g., BotChase [1]). We construct a bipartite session–URL interaction graph from standard logs and apply an inductive GNN, GraphSAGE [2], to classify session nodes. We evaluate accuracy gains and robustness to graph perturbations and temporal shift, including cold-start scoring for unseen sessions and pages.
@@
  • Non-intrusive graph formulation: We formulate bot detection on a session–URL graph from standard server logs, avoiding CAPTCHAs and client-side instrumentation.
  • Inductive GraphSAGE with lightweight features: We use GraphSAGE over session and URL attributes to score unseen session/URL nodes.
  • Robustness + deployability evaluation: We report gains over a session-feature MLP baseline and study adversarial edge perturbations and cold-start generalization.
@@ VI Conclusion
We presented a non-intrusive, graph-based bot detection framework that models session–URL interactions and applies inductive GraphSAGE for session classification. On real-world traffic with high-confidence bot labels, the refined-graph model improves over a strong session-only MLP and remains robust under mild edge perturbations and cold-start evaluation, supporting deployment as a bounded-latency backend scoring module.